Leticia gained 15 pounds in her freshman year of college. She couldn’t figure out why. She didn’t think she was eating any differently than she did at home. She is a healthy 18-year-old who is 5 feet 4 inches tall and now weights 155 lb.

Q: What is her BMI? Is it in the healthy range? Does she need to lose weight?
Answer #1

ANSWER.

Q. Calculation of BMI.

        Formula

            BMI = Weight(kg) / Height (m)2

      Here

            Weight =155 lb.

        Convert lb to kg for calculation

                1kg = 2.205 lb

          So

              Weight (kg) = 155/2.205

                                  = 70.3 kg.

             Height = 5 feet 4 inch

      Convert 5 feet 4 inch to meter for calculation.

           1 feet = 0.305 m

         So

              5 feet 4 inch = 1.62 m

    As per above formula

              BMI = 70.3 / (1.62)2

                        = 70.3 / 2.62

                        = 26.83

* Her BMI is 26.83.

* BMI between 25 to 29.9 indicate overweight.

* So she is an overweight patient and need to loose weight for a

     healthy life.